OpenStudy (quickstudent):

Can someone please check my work on transformations in exponential functions?

OpenStudy (quickstudent):

1. y = 3 * 2^x A transformation up one unit and right three units. My answer: y = 3 * (2^x - 3) + 1 the graph is in the attached file.

OpenStudy (phi):

yes, if you put the parens like this y = 3* 2^(x-3) + 1

Directrix (directrix):

This y = 3 * (2^x - 3) + 1 is y = 3*2^x -8. Compared with the given function: y = 3 * 2^x it appears that the graph dropped 8 units and did not shift to the right.

OpenStudy (quickstudent):

Ok, here's all my corrections: 1. y = 3 * 2^x A transformation up one unit and right three units. My answer: y = 3 * 2^(x - 3) + 1 the graph is in the attached file.

OpenStudy (quickstudent):

@phi

OpenStudy (phi):

yes, that is ok

OpenStudy (quickstudent):

Thanks Can you please check two more?

OpenStudy (quickstudent):

2. y = (0.5)^x A transformation down five units and left two units. My answer: y = (0.5)^(x+2) - 5 Here's the graph:

OpenStudy (phi):

yes

OpenStudy (quickstudent):

Ok:) one more 3. y = 2 * 3^x A transformation up 4 units and left 3 units. My answer: y = 2 * 3^(x + 3) + 4 Graph:

OpenStudy (phi):

yes

OpenStudy (quickstudent):

:D Thanks for all your help:)

OpenStudy (phi):

yw